Social aspects on muds

Workshop at World Mud Conference 1994

Linus Tolke

This is a summary of the discussions conducted at the social aspects workshop during the World Mud Conference 1994 in Linkoping.

The summary is made by Linus and it may not exactly describe what we talked about. This is how I remember/percieved the discussion.

Participants in this workshop were Aronsson, Elessar, MeTa, Linus, Mats, Sharilla, Swampnose, Arkon, Lythende and Mentar.

We decided to devide the discussions in two main issues:

Social life in the MUD How does it work? What rules should be obeyed?

Social life outside the mud Do the mudders have a life

The most important conclusion we arrived at is:

Mudding should be fun!

1 Social life in the MUD

1.1 Privacy and snooping

The conflict of snooping against privacy was considered.

The problem is that almost anyone can become a wizard and get the powers of snooping or in some other way seeing/logging what other players do.

These powers can easily be used to follow a players proceedings in the mud and to find what the player is missing for clues and or what he really is using in the solving of a quest, a problem.

Since it is a very simple tool it can easily be used to herass a player or it could be used to watch what a player is talking about.

Since there is a lot of freedom for wizards and because of the basic concepts of the world it is very hard to technically stop the wizards from doing some logging or watching.

Is snooping really legal? We concluded that we don’t really know exactly what the law is and anyway since muds are distributed over several countries, what country’s laws should be used?

Possible solutions to consider were suggested:

•Players must be told whenever they are or can be snooped This is really a simple solution to a complex problem. It enforces the idea that the Mud is just a game and is not supposed to be used for private conversations and could even be hindering the social parts of the Mud.

•Remove say and tell from the snooping For following the paths of a player it is normally not needed to see what he is saying or telling other players. Most private conversations are held with say and tell so by removing that possibility we remove most of the problem. This is not satisfactury of coarse since any wizard can make an object that logs or tells him what the player he is interested in is saying or telling.

•Awareness and responsibility The new wizards should be made aware of the fact that their new powers are very powerful and can be used to violate the privacy of the players.

If everybody have a suitable moral code there won’t be a problem with this.

1.2 Player killing

Player killing is probably the worst possible thing for the social life in the MUD. The person that has been killed or attacked can never trust another player again.

1.3 Quest solutions

There might be a problem with given away quest solutions. This is a sabotage against the game and it is not fun to recieve them.

1.4 Observations of the life in the MUD

Muds are a special social community. People interact in another way than they would in their normal life. This is due to the medium.

Since feelings are not shown upon the faces of the persons but has to pass your brain to be converted in a suitable action via your fingers we can make a few observations:

•You can hide your feelings The people you are interacting with do only get the feelings you want them to see. This could be very useful if you want to hide your feelings but is a big hinderance in the social interaction.

•You behave differently in the MUD How the actions or feelings you use in the MUD are chosen, like they are chosen in the real world, by the construction of the social scene. In the MUD this means they are influenced by the setting, by how the actions look on the screen and how easy it is to type them. This is not a problem as such but an interesting observation.

2 Social life outside the MUD

Do the MUDders have a life? This is a question that is very hard to answer. We just made some observations.

2.1 Muds are actually helping people develop social skills

Since MUDs create some sort of social life it can actually help people with social disabilities to learn to interact socially.

The help it can offer is a small, simple social scene where they can act without the extra stress of being watched.

It should be mentioned that the persons that really could be helped this way are rare.